# Uncertainty, von Neumann entropy and squeezing in bipartite state of two-level atoms

### v1

@@ I.3 IV GENERALIZED RELATION BETWEEN VON NEUMANN ENTROPY AND SQUEEZING PARAMETERS IN TWO-LEVEL ATOMS
We establish generalized operational measures of the von Neumann entropies of the entangled atoms. We do this by establishing generalized relations between the von Neumann entropies of the entangled atoms and spin squeezing parameter introduced by Kitagawa et. al Kitagawa ; Jian and spectroscopic squeezing parameter used in the context of Ramsey spectroscopy, introduced by Wineland et. al Wineland2 ; Jian .
Using Eqs. (9) and (43), we get
$c_{1}^{4}-c_{1}^{2}=\Delta J_{y^{\prime}}^{2}-\Delta J_{y^{\prime}}^{4}-\frac{1}{4}$ (72)
This is a quadratic equation in $c_{1}^{2}$ whose roots are
$c_{1}^{2}=\frac{1}{2}\pm\Delta J_{y^{\prime}}\sqrt{1-\Delta J_{y^{\prime}}^{2}}$ (73)
Therefore, using Eqs. (9) and (73), we get
$c_{2}^{2}=\frac{1}{2}\mp\Delta J_{y^{\prime}}\sqrt{1-\Delta J_{y^{\prime}}^{2}}$ (74)
So, using Eqs. (65), (73) and (74), we get
@@
$log_{2}\Big{[}\frac{1}{2}\mp\Delta J_{y^{\prime}}\sqrt{1-\Delta J_{y^{\prime}}^{2}}\Big{]}$
or,
@@
Here $j=\frac{N}{2}$, where $N$ is the number of atoms. In this case we have $N=2$ and hence $j=1$. Now,
$\sqrt{\frac{2}{j}}\Delta J_{\overrightarrow{n}_{\perp}}=\xi_{s}$ (77)
is the spin squeezing parameter introduced by Kitagawa et. al Kitagawa ; Jian , where $\overrightarrow{n}_{\perp}$ is in the plane perpendicular to the mean spin vector $\langle\hat{\mathbf{J}}\rangle$.
In this case the $\overrightarrow{n}_{\perp}$ are the $x^{\prime}$ and $y^{\prime}$ directions. Now we have observed from FIG. 1 that we have squeezing in the $y^{\prime}$-quadrature. So, we take $\Delta J_{\overrightarrow{n}_{\perp}}=\Delta J_{y^{\prime}}$. Therefore, in this case
$\sqrt{\frac{2}{j}}\Delta J_{y^{\prime}}=\xi_{s_{y}},$ (78)
and hence we get with $j=1$,
@@
The spectroscopic squeezing parameter $\xi_{R}$, introduced by Wineland et. al Wineland2 ; Jian is
$\xi_{R}=\sqrt{N}\frac{\Delta J_{\overrightarrow{n}_{\perp}}}{|\langle\hat{\mathbf{J}}\rangle|}.$ (80)
In this case since $N=2$, and we have squeezing in the $y^{\prime}$-quadrature, we have
@@
The Eqs. (LABEL:3.57) and (LABEL:3.60) give generalized operational measures of von Neumann entropies of the entangled atoms.
Figure 2: Plot of the spin squeezing parameter $\xi_{s_{y}}^{2}$, spectroscopic squeezing parameter $\xi_{R_{y}}^{2}$ and the von Neumann entropies $S_{a}$ and $S_{b}$ of the entangled parts with respect to the constant $c_{1}$. Here we have taken $c_{3}=c_{4}=c_{5}=c_{6}=\frac{1}{\sqrt{2}}$. $c_{1}$ has been plotted along the horizontal axis and $\xi_{s_{y}}^{2}$, $\xi_{R_{y}}^{2}$ and $S_{a}=S_{b}$ along the vertical axis. The thick line, dashed line and the dotted line represent $\xi_{s_{y}}^{2}$, $\xi_{R_{y}}^{2}$ and $S_{a}=S_{b}$ respectively.
We see from FIG. 2 that as the spin squeezing and spectroscopic squeezing of the composite state increases the von Neumann entropies of the entangled parts increase and when there is $100\%$ spin squeezing of the composite state the von Neumann entropies of the entangled parts attain their peak values. Therefore, more is the squeezing of the composite state more is the von Neumann entropy of the entangled parts. Then as the squeezing of the composite state decreases the von Neumann entropy of the entangled parts decrease. This feature, as we see, is for the general state of two two-level atoms given in Eq. (12).
@@ I.4 V CONCLUSION
We have calculated analytically the uncertainties for a general state of two entangled two-level atoms both for the composite state and the entangled parts. We observe that the von Neumann entropies of the entangled parts are always greater than that of the composite state, but for certain range of the superposing constants of the quantum state, the uncertainties of the entangled parts are less than that of the composite state. This is not in accordance with the prevailing consensus that more is the entropy, more is the uncertainty. This feature, as we see, is for the general state of two two-level atoms.
We have also established analytically the generalized relationship between the von Neumann entropies of the entangled parts and the uncertainties, spin squeezing and spectroscopic squeezing parameters, used in the context of Ramsey spectroscopy, of the composite state. This provides a generalized operational measure of the von Neumann entropies of the entangled atoms.

### v2

@@ I.3 IV GENERALIZED RELATION BETWEEN VON NEUMANN ENTROPY AND SQUEEZING PARAMETERS IN TWO-LEVEL ATOMS
We now establish generalized relationships between the von Neumann entropies of the entangled subsystems of two-level atoms and the spin and spectroscopic squeezing parameters of the composite state. This provides operational measures for the von Neumann entropies of the entangled subsystems of two-level atoms.
Using Eqs. (6) and (43), we obtain
$c_{1}^{4}-c_{1}^{2}=\Delta J_{y^{\prime}}^{2}-\Delta J_{y^{\prime}}^{4}-\frac{1}{4}.$
This is a quadratic equation in $c_{1}^{2}$, whose roots are
$c_{1}^{2}=\frac{1}{2}\pm\Delta J_{y^{\prime}}\sqrt{1-\Delta J_{y^{\prime}}^{2}}.$ (61)
Therefore, using Eqs. (6) and (61), we obtain
$c_{2}^{2}=\frac{1}{2}\mp\Delta J_{y^{\prime}}\sqrt{1-\Delta J_{y^{\prime}}^{2}}.$ (62)
Thus, using Eqs. (60), (61), and (62), we obtain
@@
$log_{2}\Big{[}\frac{1}{2}\mp\Delta J_{y^{\prime}}\sqrt{1-\Delta J_{y^{\prime}}^{2}}\Big{]},$
or equivalently,
@@
Here, $j=\frac{N}{2}$, where $N$ is the number of atoms. In this case, we have $N=2$, and hence $j=1$.
Now, using Eq. (4) in Eq. (LABEL:3.55), and taking $j=1$, we obtain
@@
This is the generalized relationship between the von Neumann entropies of the entangled subsystems of two-level atoms and the spin squeezing parameter.
To construct a general relation between the von Neumann entropies of the entangled subsystems of two-level atoms and the spectroscopic squeezing parameter, we consider Eq. (5). In this case, since $N=2$, we have
@@
Thus, we have established a generalized relationship between the von Neumann entropies of the entangled subsystems of two-level atoms and the spectroscopic squeezing parameter. The Eqs. (LABEL:3.57) and (LABEL:3.60) provide generalized operational measures of the von Neumann entropies of the entangled atoms.
In FIG. 3, we plot the spin and spectroscopic squeezing parameters of the composite state, and the von Neumann entropies of the entangled subsystems of two-level atoms, with respect to the constant $c_{1}$.
Figure 3: Plot of the spin squeezing parameter $\xi_{s_{y}}^{2}$, spectroscopic squeezing parameter $\xi_{R_{y}}^{2}$, and the von Neumann entropies $S_{a}$ and $S_{b}$ of the entangled parts with respect to the constant $c_{1}$. Here, we have taken $c_{3}=c_{4}=c_{5}=c_{6}=\frac{1}{\sqrt{2}}$. The constant $c_{1}$ is plotted along the horizontal axis, while $\xi_{s_{y}}^{2}$, $\xi_{R_{y}}^{2}$, and $S_{a}=S_{b}$ are plotted along the vertical axis. The thick line, dashed line, and dotted line represent $\xi_{s_{y}}^{2}$, $\xi_{R_{y}}^{2}$, and $S_{a}=S_{b}$, respectively.
We see from FIG. 3 that as the spin squeezing and spectroscopic squeezing of the composite state increase, the von Neumann entropies of the entangled parts also increase, and when there is $100\%$ spin squeezing of the composite state, the von Neumann entropies of the entangled parts attain their peak values. Therefore, the greater the squeezing of the composite state, the larger the von Neumann entropies of the entangled parts. Then, as the squeezing of the composite state decreases, the von Neumann entropies of the entangled parts also decrease. This feature, as we can see, applies to the general state of two two-level atoms given in Eq. (I.2).
@@ I.4 V CONCLUSION
We have analytically calculated the uncertainties for a general state of two entangled two-level atoms, both for the composite state and the entangled parts. We observe that the von Neumann entropies of the entangled parts are always greater than that of the composite state. However, for certain ranges of the superposition constants of the quantum state, the uncertainties of the entangled parts are found to be less than those of the composite state. This observation does not align with the prevailing idea that greater entropy implies greater uncertainty. As we can see, this feature holds for the general state of two two-level atoms.
We have also analytically established generalized relationships between the von Neumann entropies of the entangled parts of two-level atoms and the uncertainties, spin squeezing, and spectroscopic squeezing parameters —used in the context of Ramsey spectroscopy —of the composite state. This provides generalized operational measures of the von Neumann entropies of the entangled atoms.